BlueLaBerry:

? x + StartFraction 4 Over 9 EndFraction = three-fourths

Eiwoh2:

Hello, welcome to QuestionCove! It seems your question isn't complete. Perhaps if you show us the full question with a screenshot, we may be able to assist you. :)

Mercury:

x + (4/9) = 3/4, I believe subtract 4/9 from both sides x = (3/4) - (4/9) find the LCM of 4 and 9, convert the fractions to the same denominator, then subtract